# Stability and asymptotic behaviour of one-dimensional solutions in cylinders

### v1

@@ 1. Introduction
[…]
@@
and study it under small perturbations of $\Omega$. To well define the functional $T$ a local uniqueness result of the positive solution $u_{\Omega}$ is needed and this is ensured by considering a nondegenerate solution $u_{\Omega}$ (see Section 3 and [3, Proposition 2.1]).
@@
[…]
@@
[…]
In the paper [3] some conditions for the stability of the pair $(\Omega_{\varphi_{L}},v_{L})$ ($v_{L}$ one-dimensional nondegenerate solution) were derived, for semilinear elliptic problems with a general nonlinearity $f(u)$, while for the case of the torsion problem, i.e. $f(u)=1$, sharp results both for instability and stability were proved (see [3, Theorem 1.4]).
@@
    $\alpha_{1}(p)=-\frac{1}{2}p^{2}(1+o_{p}(1))\qquad\mbox{as $p\rightarrow+\infty$}.$
@@
    $pu_{p}^{p-1}\rightarrow\frac{\pi^{2}}{4}\qquad\text{in $C^{1}_{loc}(I)$\qquad\text{as $p\rightarrow 1$};}$
@@
[…]
@@
When stability holds we have that the cylinder with the one-dimensional solution there definite locally minimizes the energy functional with the volume constrained. This does not hold when $(\Omega_{\varphi_{L}},u_{\varphi_{L}})$ is unstable, hence there are domains close to the bounded cylinder and with the same volume which are better candidates to optimize the energy.
@@
Finally, we recall that the study of the critical pairs of the energy functional $T$ is equivalent to the study of domains for which the overdetermined problem (with an extra homogeneous Neumann condition on the relative boundary) admits a solution, see [3, Proposition 2.6].
@@ 2.1. Asymptotic behaviour as $p\rightarrow+\infty$
[…]
@@
$\alpha_{1}(p)\mu_{p}^{2}\xrightarrow[p\rightarrow\infty]{}\beta_{1}<0.$
@@
is positive and even in $I$. Furthermore, being $k^{\prime\prime}_{p}(0)=p\|u_{p}\|_{\infty}^{p-1}(\lambda_{p}\mu_{p}^{2}-1)k_{p}(0)<0$ for $p$ large enpugh and $k^{\prime}_{p}(0)=0$:
@@
and takes the form (see [9, Lemma 4.2])
@@
$\alpha_{1}(p)=-\frac{1}{2}p^{2}(1+o_{p}(1))\qquad\text{as}\quad p\rightarrow+\infty.$
@@
$\begin{cases}-\psi^{\prime\prime}_{1}=e^{W}\psi_{1}+\beta_{1}\psi_{1}\quad\mbox{ in }\mathbb{R}\\ \psi_{1}\geq 0.\end{cases}$
@@ 2.2. Asymptotic behaviour as $p\rightarrow 1$
[…]
@@ 3.1. Energy functional and energy stationary pairs
Let $\bar{\varphi}\in C^{2}(\omega)$ and let $u_{\Omega_{\bar{\varphi}}}\in W^{1,\infty}(\Omega_{\bar{\varphi}})\cap W^{2,2}(\Omega_{\bar{\varphi}})$ be a positive nondegenerate solution to (1.1), with $\Omega=\Omega_{\bar{\varphi}}$. In [3, Proposition 2.1] it has been shown that under such deformations of $\Omega_{\bar{\varphi}}$ the nondegeneracy of $u_{\Omega_{\bar{\varphi}}}$ induces a local uniqueness result for solutions of (1.1) in the deformed domains. Namely, given $v\in C^{2}(\bar{\omega})$, there exists $\delta>0$ such that for any $t\in(-\delta,\delta)$ problem (1.1) with $\Omega=\Omega_{\bar{\varphi}+tv}$ admits a unique positive solution $u_{\Omega_{\bar{\varphi}+tv}}$ in a neighborhood of $u_{\Omega_{\bar{\varphi}}}\circ\xi(t,\cdot)^{-1}$.
@@
Moreover in [3, Lemma 4.1] the first derivative of $T$ at $\bar{\varphi}$, i.e. for $t=0$, with respect to variations $v\in C^{2}(\bar{\omega})$ has been computed and it takes the following form
@@ 3.2. One-dimensional solutions
By decomposing the spectrum of $L_{u_{p,L}}=-\Delta-pu_{p,L}^{p-1}$, sufficient conditions for nondegeneracy of $u_{p,L}$ have been obtained in [3, Corollary 4.7].
@@
$\begin{cases}h^{\prime\prime}_{p,L}=(\lambda_{1}(\omega)-pu_{p,L}^{p-1})h_{p,L}\quad\mbox{ in }(0,L)\\ h_{p,L}(L)=-Lu^{\prime}_{p,L}(L)\\ h^{\prime}_{p,L}(0)=0.\end{cases}$
@@
In fact in [3] this result has been obtained in the case $L=1$, but it can be easily extended to the general case $L>0$, and more importantly in [3] these conditions are stated just as sufficient ones, but actually they hold as “if and only if”, as it is clear from the proofs of [3, Theorem 4.11] and [3, Theorem 1.5].
It is worth pointing out that assumption $\lambda_{1}(\omega)>-\alpha_{1,L}(p)$ is not only a sufficient condition to guarantee the nondegeneracy of the one-dimensional solution $u_{p,L}$ of (1.1) in $\Omega_{\varphi_{L}}$ (see Remark 3.5), but also implies the positivity of the function $h_{p,L}$, which is a key property in the proof of Theorem 3.6. We state the positivity result, obtained in [3, Proposition 4.10] via maximum principle, in a suitable form and for $L=1$.
@@
[…]
Whereas we claim that, for $p$ sufficiently large, if $\lambda\in(-\alpha_{1}(p),p\|u_{p}\|_{\infty}^{p-1})$
@@
We notice that the interval $(-\alpha_{1}(p),p\|u_{p}\|_{\infty}^{p-1})$ is not empty by (2.11).
@@
Suppose by contradiction that there exists $p_{n}\rightarrow+\infty$ and $\lambda_{n}\in(-\alpha_{1}(p_{n}),p_{n}\|u_{p_{n}}\|_{\infty}^{p_{n}-1})$ such that $\|h_{p_{n}}\|_{\infty}=h_{p_{n}}(0)$, then
$\frac{-\alpha_{1}(p_{n})}{p_{n}\|u_{p_{n}}\|_{\infty}^{p_{n}-1}}<{\lambda_{n}}\mu_{p_{n}}^{2}<1,$
@@
So, by $(3)$ and $(5)$ of Theorem 1.1, ${\lambda_{n}}\mu_{p_{n}}^{2}\rightarrow 1$, as $n\rightarrow+\infty$, and rescaling the function by
@@
$\tilde{h}^{\prime\prime}_{p_{n}}(y)=\left(\mu_{p_{n}}\lambda_{n}-\left(1+\frac{\tilde{u}_{p_{n}}(y)}{p_{n}}\right)^{p_{n}-1}\right)\tilde{h}_{p_{n}}(y)\qquad y\in\left(-\frac{1}{\mu_{p_{n}}},\frac{1}{\mu_{p_{n}}}\right).$
@@
$\begin{cases}\tilde{H}^{\prime\prime}=(1-e^{W})\tilde{H}\qquad\text{in $\mathbb{R}$}\\ \tilde{H}(0)=1,\tilde{H}^{\prime}(0)=0.\end{cases}$
Being $\tilde{H}>0$ and
$1-e^{W(y)}=\frac{(1-e^{\sqrt{2|y|}})^{2}}{(1+e^{\sqrt{2|y|}})^{2}}>0\quad\mbox{ if }y\neq 0,$
we get that $\tilde{H}>1$ in $(-\varepsilon,\varepsilon)\setminus\{0\}$ which is impossible considering that $\|\tilde{h}_{p_{n}}\|_{\infty}=1$. This proves (3.7) and concludes the proof. ∎
@@
[…]
@@
Let $\eta\in(0,1)$ be fixed, then by by $(3)$ of Theorem 1.2 we have $pu_{p}^{p-1}\rightrightarrows\frac{\pi^{2}}{4}$ in $\eta,\eta]$. Furthermore

### v2

@@ 1. Introduction
where $p>1$. By standard variational methods it is easy to see that a positive weak solution of (1.1) exists in the Sobolev space $H^{1}_{0}(\Omega\cup\Gamma_{1,\Omega})$, which is the space of functions in $H^{1}(\Omega)$ whose trace vanishes on the relative boundary $\Gamma_{\Omega}$, at least for some values of $p$.
Let $u_{\Omega}$ be a positive solution of (1.1), then its energy is defined as:
@@
and study it under small perturbations of $\Omega$. To well define the functional $T$ a local uniqueness result of the positive solution $u_{\Omega}$ is needed and this is ensured by considering a nondegenerate solution $u_{\Omega}$ (see Section 3 and [4, Proposition 2.1]).
@@
In this domain there exists a one-dimensional solution of (1.1) which is easily obtained by extending to $\Omega_{\varphi_{L}}$ a positive solution $u_{p,L}$ of the following ODE problem:
@@
In view of the simple geometry of $\Omega_{\varphi_{L}}$ and of the corresponding one-dimensional solution $u_{p,L}$, we may ask wether $\Omega_{\varphi_{L}}$ is or not a good candidate to locally minimize the energy functional $T$. This question is addressed in [4] (see also [2]) by differentiating $T$ with respect to variation of $\Omega$ which leave the measure invariant and studying the stability/instability of the pair $(\Omega_{\varphi_{L}},u_{p,L})$ as a critical point of $T$ (see Section 3 for the definition and the setting). Indeed, when $u_{p,L}$ is a nondegenerate solution of (1.1) (see Proposition 3.4) we may consider the energy functional $T({\Omega_{\varphi}})$ for domains $\Omega_{\varphi}$ close to $\Omega_{\varphi_{L}}$.
In the paper [4] some conditions for the stability of the pair $(\Omega_{\varphi_{L}},v_{L})$ ($v_{L}$ one-dimensional nondegenerate solution) were derived, for semilinear elliptic problems with a general nonlinearity $f(u)$, while for the case of the torsion problem, i.e. $f(u)=1$, sharp results both for instability and stability were proved (see [4, Theorem 1.4]). A related result is obtained in [3] by showing a bifurcation result from 1-dimensional solutions of semilinear elliptic problems in bounded cylinders.
@@
    $\alpha_{1}(p)=-{\color[rgb]{1,0,0}\frac{1}{4}}p^{2}(1+o_{p}(1))\qquad\mbox{as $p\rightarrow+\infty$}.$
@@
    $pu_{p}^{p-1}\rightarrow\frac{\pi^{2}}{4}\qquad\text{in $C^{1}_{loc}(I)$\hskip 20.44434pt\text{as $p\rightarrow 1$};}$
@@
Let $\omega\subset\mathbb{R}^{N-1}$ be a smooth bounded domain, let $\lambda_{1}(\omega)$ be the first non-trivial eigenvalue of the Laplacian in $\omega$ (i.e. in (N1)(N-1) coordinates) with Neumann boundary condition, let $u_{p,L}$ be the positive one-dimensional solution to (1.1) in $\Omega_{\varphi_{L}}$ defined in (1.6) and let $\alpha_{1}(p)$ be the first eigenvalue of (1.8) with Dirichlet boundary conditions in I. Then, for any $\gamma>\frac{1}{2}$, if $p$ is sufficiently large and $L>\frac{\gamma\,p}{\sqrt{\lambda_{1}(\omega)}}$, the pair $(\Omega_{\varphi_{L}},u_{p,L})$ is a stable energy-stationary pair (see Definition 3.2).
@@
When stability holds we have that the cylinder with the one-dimensional solution there definite locally minimizes the energy functional with the volume constrained. This does not hold when $(\Omega_{\varphi_{L}},u_{\varphi_{L}})$ is unst0able, hence there are domains close to the bounded cylinder and with the same volume which are better candidates to optimize the energy.
@@
Finally, we recall that the study of the critical pairs of the energy functional $T$ is equivalent to the study of domains for which the overdetermined problem (with an extra homogeneous Neumann condition on the relative boundary) admits a solution, see [4, Proposition 2.6].
@@ 2.1. Asymptotic behaviour as $p\rightarrow+\infty$
Being $p\leq\tilde{u}_{p}\leq 0$, we deduce that $0\leq|\tilde{u}^{\prime\prime}_{p}|\leq 1$, thus in any compact subset of $\mathbb{R}$ $\tilde{u}^{\prime}_{p}$ is uniformly bounded and then Ascoli-Arzelà theorem implies that up to a subsequence $\tilde{u}_{p}\rightarrow W$ in $C^{1}_{loc}(\mathbb{R})$, where $W$ solves (1.12). It is easy to see that the solution of (1.12) takes the form
@@
$\alpha_{1}(p)\mu_{p}^{2}\xrightarrow[p\rightarrow\infty]{}\beta_{1}<0.$ (2.9)
@@
is positive and even in $I$. Furthermore, being $k^{\prime\prime}_{p}(0)=p\|u_{p}\|_{\infty}^{p-1}(\lambda_{p}\mu_{p}^{2}-1)k_{p}(0)<0$ for $p$ large enough and $k^{\prime}_{p}(0)=0$:
@@
and takes the form (see [10, Lemma 4.2])
@@
$\alpha_{1}(p)={\color[rgb]{1,0,0}-\frac{1}{4}}p^{2}(1+o_{p}(1))\qquad\text{as}\quad p\rightarrow+\infty.$
@@
$\begin{cases}-\psi^{\prime\prime}_{1}=e^{W}\psi_{1}+\beta_{1}\psi_{1}\quad\mbox{ in }\mathbb{R}\\ \psi_{1}\geq 0,\end{cases}$
where $\beta_{1}$ is the constant introduced in (2.9).
@@ 2.2. Asymptotic behaviour as $p\rightarrow 1$
Now, multiplying (2.21) by $\varphi_{1}$, (2.22) by $\bar{u}_{p}$ and integrating both equations in $I$, we obtain
@@ 3.1. Energy functional and energy stationary pairs
Let $\bar{\varphi}\in C^{2}(\omega)$ and let $u_{\Omega_{\bar{\varphi}}}\in W^{1,\infty}(\Omega_{\bar{\varphi}})\cap W^{2,2}(\Omega_{\bar{\varphi}})$ be a positive nondegenerate solution to (1.1), with $\Omega=\Omega_{\bar{\varphi}}$. In [4, Proposition 2.1] it has been shown that under such deformations of $\Omega_{\bar{\varphi}}$ the nondegeneracy of $u_{\Omega_{\bar{\varphi}}}$ induces a local uniqueness result for solutions of (1.1) in the deformed domains. Namely, given $v\in C^{2}(\bar{\omega})$, there exists $\delta>0$ such that for any $t\in(-\delta,\delta)$ problem (1.1) with $\Omega=\Omega_{\bar{\varphi}+tv}$ admits a unique positive solution $u_{\Omega_{\bar{\varphi}+tv}}$ in a neighborhood of $u_{\Omega_{\bar{\varphi}}}\circ\xi(t,\cdot)^{-1}$.
@@
Moreover in [4, Lemma 4.1] the first derivative of $T$ at $\bar{\varphi}$, i.e. for $t=0$, with respect to variations $v\in C^{2}(\bar{\omega})$ has been computed and it takes the following form
@@ 3.2. One-dimensional solutions
By decomposing the spectrum of $L_{u_{p,L}}=-\Delta-pu_{p,L}^{p-1}$, sufficient conditions for nondegeneracy of $u_{p,L}$ have been obtained in [4, Corollary 4.7].
@@
$\begin{cases}h^{\prime\prime}_{p,L}=(\lambda_{1}(\omega)-pu_{p,L}^{p-1})h_{p,L}\quad\mbox{ in }(0,L)\\ h_{p,L}(L)=-Lu^{\prime}_{p,L}(L)\\ h^{\prime}_{p,L}(0)=0.\end{cases}$ (3.3)
@@
In fact in [4] this result has been obtained in the case $L=1$, but it can be easily extended to the general case $L>0$, and more importantly in [4] these conditions are stated just as sufficient ones, but actually they hold as “if and only if”, as it is clear from the proofs of [4, Theorem 4.11] and [4, Theorem 1.5].
It is worth pointing out that assumption $\lambda_{1}(\omega)>-\alpha_{1,L}(p)$ is not only a sufficient condition to guarantee the nondegeneracy of the one-dimensional solution $u_{p,L}$ of (1.1) in $\Omega_{\varphi_{L}}$ (see Remark 3.5), but also implies the positivity of the function $h_{p,L}$, which is a key property in the proof of Theorem 3.6. We state the positivity result, obtained in [4, Proposition 4.10] via maximum principle, in a suitable form and for $L=1$.
@@
If $\lambda\geq p\|u_{p}\|_{\infty}^{p-1}$ then, for any $p>1$, by (3.4) we have $h^{\prime\prime}_{p}>0$ in $(0,1)$, which implies $h^{\prime}_{p}(1)>h^{\prime}_{p}(0)=0$. This case was already covered by [4].
Whereas we claim that, given $\gamma>\frac{1}{2}$, for $p$ sufficiently large, if $\lambda\in[{\color[rgb]{1,0,0}\gamma^{2}p^{2}},p\|u_{p}\|_{\infty}^{p-1})$
@@
We notice that the interval ${\color[rgb]{1,0,0}[\gamma^{2}p^{2}},p\|u_{p}\|_{\infty}^{p-1})$ is not empty by Proposition 2.7.
@@
Suppose by contradiction that there exists $p_{n}\rightarrow+\infty$ and $\lambda_{n}\in{\color[rgb]{1,0,0}[\gamma^{2}p_{n}^{2}},p_{n}\|u_{p_{n}}\|_{\infty}^{p_{n}-1})$ such that $\|h_{p_{n}}\|_{\infty}=h_{p_{n}}(0)$, then
$\frac{\gamma^{2}}{p_{n}\|u_{p_{n}}\|_{\infty}^{p_{n}-1}}<\mu_{p_{n}}^{2}{\lambda_{n}}<1,$
@@
So, by $(3)$ of Theorem 1.1, up to a subsequence, ${\lambda_{n}}\mu_{p_{n}}^{2}\rightarrow{\color[rgb]{1,0,0}\eta\in[2\gamma^{2},1)\subset(\frac{1}{2},1)}$, as $n\rightarrow+\infty$, and rescaling the function by
@@
$\begin{cases}\tilde{h}^{\prime\prime}_{p_{n}}(y)=\left(\mu_{p_{n}}^{2}\lambda_{n}-\left(1+\frac{\tilde{u}_{p_{n}}(y)}{p_{n}}\right)^{p_{n}-1}\right)\tilde{h}_{p_{n}}(y)\qquad y\in\left(-\frac{1}{\mu_{p_{n}}},\frac{1}{\mu_{p_{n}}}\right)\\ \tilde{h}_{p_{n}}(0)=1\\ \tilde{h}^{\prime}_{p_{n}}(0)=0\\ \tilde{h}_{p_{n}}>0.\end{cases}$
@@
$\begin{cases}-\tilde{H}^{\prime\prime}=e^{W}\tilde{H}-{\color[rgb]{1,0,0}\eta}\tilde{H}\qquad\text{in $\mathbb{R}$}\\ \tilde{H}(0)=1,\tilde{H}^{\prime}(0)=0\\ \|\tilde{H}\|_{\infty}=1.\end{cases}$
Reasoning exactly as in the last part of the proof of Theorem 2.11 we conclude that $\eta=\frac{1}{2}$, which is a contradiction against $\eta\in(\frac{1}{2},1]$.
@@
If $\lambda>\frac{\pi^{2}}{4}$, then by (1.13), for $p$ sufficiently close to $1$, $\lambda>pu_{p}^{p-1}$ in $[0,1]$ and so, by (3.4), for $p$ sufficiently close to $1$, $h^{\prime\prime}_{p}(t)>0$ for all $t\in[0,1]$, which implies $h^{\prime}_{p}(1)>h^{\prime}_{p}(0)=0$.
@@
Let $\eta\in(0,1)$ be fixed, then by $(3)$ of Theorem 1.2 we have $pu_{p}^{p-1}\rightrightarrows\frac{\pi^{2}}{4}$ in $\eta,\eta]$. Furthermore